OpenStudy (anonymous):

Solve the absolute value equation. |3x| = 12

OpenStudy (anonymous):

A. {4, 4} B. {4, –4} C. {–4, –4} D. no solution

OpenStudy (uri):

|3x| = 12 3x=12 x=12/3 x=4 |3x| = 12 3x=-12 x=-12/3 x=-4

OpenStudy (anonymous):

maybe it's 2,3

OpenStudy (anonymous):

aye uri , finally caught on to the system , guess you want medals ? XP

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go away ^

OpenStudy (uri):

No i just wanted to solve it.

OpenStudy (anonymous):

oh lol